QUESTIONS 20, 21 AND 22 ARE BASED ON THE FOLLOWING INFORMATION.

Suppose A Check conducted further research using a sample size of 100. The number of
times the AI algorithm is successful at detecting fake news is normally distributed with a different
mean of 900 and the same standard deviation of 100. Let X be the number of times the algorithm is successful at detecting fake news.

Question 20

What is the standard error of the mean?

(1) 900
(2) 100
(3) 10
(4) 1
(5) 0

Question 21

What is the mean of the sampling distribution of the mean?

(1) 900
(2) 100
(3) 10
(4) 1
(5) 0

Question 22

P(X ≤ 920) =?

(1) 2.0000
(2) 0.9772
(3) 0.5793
(4) 0.2000
(5) 0.0228
1
Expert's answer
2020-05-06T20:10:49-0400

20)(3) 10

e=sd/√n=100/10=10

21) (1) 900

mean of the sampling distribution is 900

22) (3) 0.5793

P(x<920)

z=(920-900)/100=0.2

P(z<0.2)=0.5793